Given the graphs of f(x) = x - 7 and g(x) = -5x - 1, what is the solution to the equation f(x) = g(x)
Answer by dkppathak(439)

Given the graphs of f(x) = x - 7 and g(x) = -5x - 1, what is the solution to the equation f(x) = g(x)
Solution
f(x) = g(x)
x-7=-5x-1
x+5x=-1+7
6x=6
x=1